OpenStudy (anonymous):

Jack was 11 years older than priscilla. Together their ages totaled 161 years. What were their ages ? HELP PLEASEE

OpenStudy (imstuck):

Since Jack's age is based on Priscilla's age, let's call her age x, and since Jack is 11 years older, his age is x + 11. Together their ages add up to 161, so Jack's age, x + 11, plus her age, x, = 161. x + (x + 11) = 161

OpenStudy (imstuck):

See that? See how it works?

OpenStudy (anonymous):

everyone wants to do algebra, no one wants to think i guess 70 and 81, but that is wrong because \(70+81=151\) so i am too low try higher

OpenStudy (yolomcswagginsggg):

answer 150

OpenStudy (yolomcswagginsggg):

answer 150!!!!!!!!!!

OpenStudy (anonymous):

70 AND 91 ?

OpenStudy (yolomcswagginsggg):

1 5 0 fcking 150